Whilst shopping, the probability that Caroline buys fruit is 0.6.
The probability that Caroline buys a CD is 0.2.
Buying fruit and buying a CD are independent of each other.
Work out the probability that she buys fruit, a CD or both
1
Expert's answer
2019-11-19T13:04:09-0500

Probability of buying fruit "P(f)=0.6"

Probability of buying a CD "P(c)=0.2"

Since the two actions are independent,

Probability of buying both "P(fc)=P(f)\u00d7P(c)"

Thus, "P(fc)=0.6\u00d70.2=0.12"

Probability of buying fruit, a CD or both is "P(f)+P(c)-P(fc)" "=0.6+0.2-0.12=0.68"

Answer: 0.68.